OpenStudy (anonymous):

what is the difference between dv and Δv? here is my problem: a sphere increases in diameter from 5'' to 5.1'' a)calculate dv b)calculate Δv c)calculate the error

OpenStudy (anonymous):

dv is used in calculus to represent an infitescimal. An infiniteschimal can be thought of as a change (in this case a change in v) that is almost zero. Actually, it is the closest "number" to zero without being zero. On the other hand, Δv is simply the difference (final volume - initial volume). Does this help?

OpenStudy (anonymous):

ok.. so Δv would be .1 but im still confused with dv? how would you calculate that.

OpenStudy (anonymous):

I don't think that's what it is asking. I think it is talking about the change in volume of the sphere when the diameter is changed by 0.1 and the infintescimal change in volume at d=5.

OpenStudy (anonymous):

im really confused... lol.

OpenStudy (anonymous):

Δv = V₂ - V₁ dv is derivative of V ( instantaneous volume rate)

OpenStudy (anonymous):

*rate of change

OpenStudy (anonymous):

ok.. so does that mean that dv is just 0?

OpenStudy (anonymous):

No, you need to take derivative to get dv !

OpenStudy (anonymous):

but i thought the derivative of 5 is 0

OpenStudy (anonymous):

No, take derivative of V ( Volume of Sphere formula)

OpenStudy (anonymous):

oooh ok so dv is \[4\pi r^2\] and then Δv is just .1?

OpenStudy (anonymous):

Oh, you don't know how to find dv, don't you! -> dv = 4π r² dr

OpenStudy (anonymous):

@zepdrix lol.

zepdrix (zepdrix):

sandy potato \:D/

OpenStudy (anonymous):

♥ help me.

zepdrix (zepdrix):

Calculating error? :U Oh crapsicle... I don't remember doing this.. .hmmm.

zepdrix (zepdrix):

sec thinking c:

OpenStudy (anonymous):

@zepdrix Percentage error: dv/ V

zepdrix (zepdrix):

Hmm let's start with ΔV first. So they told us the change in DIAMETER. We're going to write those diameters in terms of Radii, and then calculate the 2 volumes that they give us. Then, as Phyll stated, we'll be able to calculate the AVERAGE change in volume (ΔV) By taking the difference of the volumes! ΔV=V₂ - V₁

zepdrix (zepdrix):

For the first volume, we'll call V₁, it has a diameter of 5" which will give us a radius of 2.5" right? :o

zepdrix (zepdrix):

If we plug that into our equation for a sphere, we get that V₁ is,\[\huge V_1=\frac{4}{3}\pi (2.5)^3 \approx 65.4498\] Still wif me? :3

OpenStudy (anonymous):

i get it (:

zepdrix (zepdrix):

So V₂ will be the same process. A diameter of 5.1" will give us a radius of 2.55" I think... So for V₂, that gives us...\[\huge V_2=\frac{4}{3}\pi(2.55)^3 \approx 69.4559\]

OpenStudy (anonymous):

Δv = V₂ - V₁ = (4π/3) ( 2.55 - 2.5)³

zepdrix (zepdrix):

Pshhhh look at you getting all fancy with your factoring :D hah

zepdrix (zepdrix):

how do you do those cute little subscripts and superscripts in normal text like that?? D:

OpenStudy (anonymous):

ZEP ITS CALLED COPY AND PASTE FROM GOOGLE

OpenStudy (anonymous):

or that.

zepdrix (zepdrix):

Oh goodness :P bunch of noobs :3

zepdrix (zepdrix):

Ok ok ok ok so to get ΔV, we simply take the difference of those 2 volumes, what'd you get? :O

OpenStudy (anonymous):

@ksandoval Pleaze read our explanation :)

OpenStudy (anonymous):

oh wait. nevermind

zepdrix (zepdrix):

Nooooo :O Δd, the difference in diameters is 5.1 - 5 = 0.1 Δr, the difference in radii is 2.55 - 2.5 = 0.05 ΔV, the difference in volumes is 69.4559 - 65.4498 = ?

OpenStudy (anonymous):

and calculate with us, would you :)

OpenStudy (anonymous):

4.0061 x.o

zepdrix (zepdrix):

cool, so the AVERAGE rate of change of volume over whatever the time interval was, is about 4 cubic inches! c: cool. ΔV, average rate of change dV, instantaneous rate of change Ok we got the first part done it looks like! :D

OpenStudy (anonymous):

well wait, so is dv just the derivative of the formula for the volume of a sphere?

OpenStudy (anonymous):

-.- ok.

zepdrix (zepdrix):

lolol simmer down phyll XD that happens sometimes. students need to hear things multiple times before it actually sinks in :D don't get frustrated heh.

OpenStudy (anonymous):

well i was just confused because i thought for the error you need to subtract dv from Δv, but dv is a formula..

zepdrix (zepdrix):

Hmm I'm a little confused about the dV part myself. Do they want you to calculate the derivative at 5 or at 5.1? Maybe @Chlorophyll knows? D:

OpenStudy (anonymous):

dv with r = 2.5 and r' = .05

zepdrix (zepdrix):

Oh oh oh ok that makes sense :D

OpenStudy (anonymous):

-> dv = 4π r² dr

OpenStudy (anonymous):

ok. sooo error is .0791?

zepdrix (zepdrix):

Oh sorry :O hmm

zepdrix (zepdrix):

What was dv? 3.93 or something like that? :O

zepdrix (zepdrix):

dV=4π r² dr --> dV=4π (2.5)² (.05) --> dV=3.92699

zepdrix (zepdrix):

\[\large \frac{dV}{V}=\frac{3.92699}{V_1}\]I assume we're using V1 for the volume? Like we did with the radius earlier? :o

zepdrix (zepdrix):

I'm coming up with 0.06 for the error. hmm

zepdrix (zepdrix):

Hmm, I'm not enjoying these problems Sand... -_- do some more fun math lol